On January 1, 2011 Murphy Co. purchased a truck that cost $36,000. The truck had an expected useful life of 10 years and a $4,000 salvage value. The amount of depreciation expense recognized in 2012 assuming that Murphy uses the double declining-balance method is:
Answer #1

Depreciation rate = 200/Useful life = 200/10 = 20%

Depreciation for 2011 = 36,000 * 20% = 7,200

Depreciation for 2012 = (36,000-7,200)*20%

= 5760

Comment if you face any issues